# Bayesian Optimization in Language Space: An Eval-Efficient AI Self-Improvement Framework

### v2

@@ 1.5 Theoretical Guarantees and Empirical Evaluation
In sum, our paper makes four main contributions to the literature on AI, optimization, and data-driven decision making. First, from a conceptual perspective, we identify evaluation efficiency as the key objective to optimize in self-improving AI design in societal and business systems. Second, we propose T-BoN BO, a purely text-based self-improving AI framework that combines textual-gradient search with Best-of-$N$ selection to automate and optimize business and marketing decisions (e.g., ad creatives, product descriptions). Third, we provide theory showing that Best-of-$N$ over locally sampled textual edits selects, in probability, ascent directions of the UCB acquisition with exploration weight $\beta=\Theta(\sqrt{\ln N})$; thus, we prove that T-BoN BO emulates gradient-based UCB and inherits evaluation-efficiency guarantees. Finally, we demonstrate empirical gains on automated ad-alignment across eight scenarios: T-BoN BO improves faster and attains higher final scores compared to strong baselines, including Best-of-64 and GEPA, and remains effective both in settings with and without contextual information.
The rest of the paper is organized as follows. Section 2 discusses the related literature. Section 3 formalizes the AI system and its objective, along with self-improving algorithms. Section  4 describes the concept of Best-of-N gradient and proves that it theoretically emulates the gradient of the UCB acquisition function. Section 5 introduces the main self-improving AI algorithm proposed in this paper, T-BoN BO. Sections 6 and 7 empirically demonstrate T-BoN BO’s effectiveness on ad-alignment tasks versus strong baselines. Finally, Section 8 concludes with a discussion of implications and limitations.
@@ 2 Related Literature
The literature on self-improving AI has often been motivated by an observation: the progress in AI based only on human data is hitting limits, so AI must create and learn from its own data. In other words, powerful AI “should have their own stream of experience that progresses, like humans, over a long time-scale” (silver2025welcome). That is, we need to design the self-improving AI that iteratively 1) generate data of outcomes, evaluate outcomes with grounded signals, and 2) self-tune the three knobs –models, tools, and prompts– in the system to achieve long-term objectives under real-world feedback (wang2025maestro). Here, tuning models is about altering language models’ input-output correspondences (huang-etal-2023-large; lee2023rlaif; acikgoz2025self); tuning tools is about specifying the tools the AI has access to, including memory tools (hou2025model; ouyang2025reasoningbank; qiu2025alita); tuning prompts is about engineering prompts in the model’s context window, including system messages, task instructions, constraints, schemas, and few-shot exemplars (zhou2022large; khattab2023dspy; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro). In this paper, we focus on studying self-improving systems based on tuning prompts.
@@
Our paper also relates to the literature on relating LLMs to Bayesian Optimization (BO). A series of papers focuses on applying LLM to enhance BO for numerical optimization problems. liularge and cisse2025language focus on leveraging LLMs’ contextual understanding and few-shot learning capabilities to improve BO’s efficiency, enabling better warm-starting and surrogate modeling; agliettifunbo uses an LLM to discover new acquisition functions for Bayesian optimization. The LLM writes candidate acquisition function formulas as code, which are evaluated on various optimization problems. Conversely, singh2025mechanistic applies BO to enhance LLM’s internal behavior to boost the model’s zero-shot and few-shot performance. In contrast to these papers, our paper focuses on considering the prompt optimization problem that arises in formulating LLM-based self-improving AI system design as a Bayesian Optimization problem. In line with this focus, schneider2025hyperbandbased and kong2025metapromptoptimizationllmbasedsequential formulate prompt optimization among the fixed set of candidate prompts as Bayesian optimization. While schneider2025hyperbandbased utilizes ideas from adversarial bandits, kong2025metapromptoptimizationllmbasedsequential feeds each prompt text into an embedding model, and the GP surrogate operates on those fixed embedding features to guide selection. In contrast to these approaches, our work focuses on self-improving AI that actively generate new candidate prompts as the process proceeds, rather than pre-specifying a fixed set of candidate prompts.
@@
In terms of method evaluation benchmarks, our paper relates to the methodological literature on ad optimization in digital marketing. A series of papers focused on the efficient execution of ad evaluation, i.e., iteratively improving ad choice efficiently given a fixed set of candidate ads. schwartz2017customer adopts Thompson Sampling with a hierarchical general linear model for a real display advertising campaign for a financial-services firm, which purchased display advertising across 59 websites. geng2020online partitions JD.com’s platform users into disjoint sub-populations and applies a contextual Thompson sampling algorithm to maximize the expected advertiser payoff. aramayo2023multiarmed developed a contextual Thompson sampling algorithm to dynamically determine the list of house ads to display on the homepage of an electronic retailer to maximize the accumulated click-through rate of the ads. ba2022advertising used parametric Thompson Sampling, where a parametric structure links features (media, audience attributes) to conversion likelihood to optimally allocate ad-media and target-audience combinations in high-dimensional settings with low success rates. The main difference between bandit-based papers and ours is the existence of a pre-defined set of candidate arms. Self-improving AI focuses on creating new solutions by analyzing previous results. On the other hand, bandit methods aim to balance exploration of a predefined set of candidate ads optimally. More broadly, there is also a difference in problem scope. Self-improving AI iterates among three steps: generation, evaluation, and analysis. In contrast, bandit methods focus on efficiently executing the evaluation step. Indeed, an AI self-improvement framework’s evaluation step may employ a bandit method to improve efficiency further.
3 Setup and Background
3.1 The optimization objective.
@@ 3 Setup and Background
Consider a system designer’s decision-making problem, where she chooses a prompt $\Pi$ for the LLM-based AI system $\Phi$. [^3] LLM-based AI systems can also be modified by tuning the LLM’s internal parameters using LLM fine-tuning methods such as GRPO (liu2024deepseek). In this paper, we focus on prompt optimization, which fixes those internal parameters throughout. We call $\Phi(\Pi)$ the ‘system outcome’ of $\Phi$ corresponding to a prompt $\Pi$. In a digital marketing setting, $\Phi$ is an image generation AI, and $\Phi(\Pi)$ is the ad generated by the prompt $\Pi$. As a system, $\Phi(\Pi)$ has inputs and corresponding outputs. Specifically, $\Phi(\Pi)$ maps the input space $\mathcal{X}$ to a distribution over an output space $\mathcal{Y}$. For each $x\in\mathcal{X}$, we denote by $y(\Phi(\Pi),x)$ the random $\mathcal{Y}$-valued vector output of the system at input $x$. In a digital marketing setting, input $x\in\mathcal{X}$ is a platform user, and a realization of $y(\Phi(\Pi),x)$ represents that user’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\Pi)$.
The distribution of inputs over $\mathcal{X}$ may depend on the prompts set $\Pi$. For example, in a digital marketing setting, a recommender system on the ad platform determines which users are shown the ad $\Phi(\Pi)$. To capture such dependences, we denote the input distribution over $\mathcal{X}$ by $\mathcal{D}_{\mathcal{X}}(\Pi)$.
@@
To evaluate the output, we use a scoring rule: a function $r:\mathcal{Y}\to[0,1]$ that assigns a scalar effectiveness score to each $y\in\mathcal{Y}$. In the digital marketing example, $r$ is a marketer-defined function that maps observed user responses (clicks, conversions, spend) to a normalized scalar effectiveness score.
Our optimization objective is to discover the prompt $\Pi^{*}$ that maximizes the objective function
@@
Note that $J$ cannot be optimized by directly computing gradients, because the input distribution $\mathcal{D}_{\mathcal{X}}(\Pi)$ can shift with $\Pi$; even when $J(\Pi)$ is differentiable with respect to $\Pi$ and $y$’s randomness is not dependent on $\Pi$, we have:
@@
where $d_{\Pi}$ is the density of $\mathcal{D}_{\mathcal{X}}(\Pi)$. Since $d_{\Pi}$ and $\nabla_{\Pi}\log d_{\Pi}(x)$ are unobservable, we cannot compute gradients from observed rewards. This motivates the use of an external LLM that can analyze prior data and propose directions for improvement without explicitly computing gradients, leading to self-improving AI algorithms we describe below that iteratively optimize for $J$.
3.2 Self-improving AI algorithm
Given a scoring rule $r$, a self-improving AI algorithm $\mathcal{A}$ optimizes $J(\Pi):=\mathbb{E}[r(y(\Phi(\Pi),x))]$ through iteratively analyzing the previous history, proposing a new prompt, and observing its noisy score. Formally, starting from the initial prompt $\Pi_{0}$, a self-improving AI algorithm proceeds in discrete optimization iterations, indexed by $t=1,2,\ldots$; at the beginning of iteration $t$, the algorithm has access to the history of past prompts and their observed scores, then selects a new prompt $\Pi_{t}$, deploys the corresponding system outcome $\Phi(\Pi_{t})$, and receives a score feedback on its performance. We define the score $s_{t}$ of prompts $\Pi_{t}$ as the empirical average reward at $t$, i.e.,
@@ 3.2 Self-improving AI algorithm
where $L$ denotes the number of input samples $x_{t}^{(l)}$ we evaluate for each $t$. Note that we compare the performance of algorithms in terms of the progress of $s_{t}$ across iterations $t$.
@@
Fix a system $\Phi$ and a scoring rule $r$. For each iteration $t\geq 1$, let the history of past evaluations be
$H_{t-1}\;:=\;\{(\Pi_{\tau},c_{\tau},s_{\tau})\}_{\tau=0}^{t-1}$ (4)
where $s_{\tau}$ is the empirically evaluated score of $\Pi_{\tau}$, defined as
@@
$\Pi_{t}=\mathcal{A}_{t}(H_{t-1})$
The goal of $\mathcal{A}$ is to choose the sequence $(\Pi_{t})_{t\geq 0}$ so as to optimize the objective $J(\Pi)$, equivalently to drive the sequence of empirical scores $(s_{t})_{t\geq 1}$ upward over iterations in expectation.
@@
    Analysis: The critic LLM analyzes the previous history $H_{t-1}$.
@@
    Generation: The critic LLM proposes the next prompts $\Pi_{t}$ based on the analysis.

### v3

@@ 1.4 Contributions
In sum, our paper makes four main contributions to the literature on AI, optimization, and data-driven decision making. First, from a conceptual perspective, we identify evaluation efficiency as the key objective to optimize in self-improving AI design in societal and business systems. Second, we propose TextBO, a purely text-based self-improving AI framework that combines textual-gradient search with Best-of-N selection to automate and optimize business and marketing decisions (e.g., ad creatives, product descriptions). TextBO is simple and easy to implement: it requires no explicit uncertainty modeling, embeddings, or Gaussian process surrogates. Third, we provide theory showing that Best-of-N over locally sampled textual edits selects, in probability, ascent directions of the UCB acquisition with exploration weight $\beta=\Theta(\sqrt{\ln N})$; thus, we prove that TextBO emulates gradient-based UCB and inherits evaluation-efficiency guarantees. Finally, we demonstrate empirical gains using a wide variety of numerical experiments, including automated ad-alignment across eight scenarios as well as the standard agentic AI benchmarks used in the self-improving AI literature – throughout, TextBO improves faster and attains higher final scores compared to strong baselines, and remains effective both in settings with and without contextual information.
@@ 2 Related Literature
The literature on self-improving AI has often been motivated by an important observation: the progress in AI based only on human data is hitting limits, so AI must create and learn from its own data. In other words, powerful AI “should have their own stream of experience that progresses, like humans, over a long time-scale” (silver2025welcome). That is, we need to design the self-improving AI that iteratively 1) generate data of outcomes, evaluate outcomes with grounded signals, and 2) self-tune the three knobs – models, tools, and prompts – in the system to achieve long-term objectives under real-world feedback (wang2025maestro). Here, tuning models is about altering language models’ input-output correspondences (huang-etal-2023-large; lee2023rlaif; acikgoz2025self); tuning tools is about specifying the tools the AI has access to, including memory tools (hou2025model; ouyang2025reasoningbank; qiu2025alita); tuning prompts is about engineering prompts in the model’s context window, including system messages, task instructions, constraints, schema, and few-shot exemplars (zhou2022large; khattab2023dspy; pryzant2023automatic; yang2024largelanguagemodelsoptimizers; textgrad; agrawal2025gepa; zhang2025agentic; wang2025maestro).
In particular, iterative prompt-optimization based self-improving AI have become increasingly popular. In a recent paper, agrawal2025gepa demonstrate their method, GEPA (Genetic-Pareto) is able to outperform state-of-the-art model fine-tuning techniques such as GRPO (liu2024deepseek) under a fixed budget of LLM generation calls across a range of challenging tasks including multi-document QA, constrained instruction following, retrieval-augmented fact verification, and privacy-related objectives. At a high-level, both GEPA and other prompt-based self-improvement (agrawal2025gepa; zhang2025agentic; wang2025maestro) is specified as procedural heuristic search: maintain a pool of prompts, generate variants via LLM-driven mutation/crossover/reflection, score candidates on held-out data or rollouts, retain top performers, and iterate (10.5555/3692070.3692611; guo2024connecting; agrawal2025gepa; better_together). This mirrors classical heuristic and evolutionary optimization, where performance is assessed as a function of generation efficiency rather than evaluation efficiency (nguyen2016understanding; ding2023quality; hu2024automated; liu2024evolution). Yet, as discussed earlier, in many real-world deployments, the scarce resource is not candidate generation but evaluation. Thus, in this paper, we focus on developing an iterative prompt-optimization framework that treat evaluation calls as the budget to allocate and optimize.
@@
Our paper also relates to the literature on relating LLMs to Bayesian Optimization (BO). A series of papers focuses on applying LLMs to enhance BO for numerical optimization problems. liularge and cisse2025language focus on leveraging LLMs’ contextual understanding and few-shot learning capabilities to improve BO’s efficiency, enabling better warm-starting and surrogate modeling; agliettifunbo uses an LLM to discover new acquisition functions for Bayesian optimization. The LLM writes candidate acquisition function formulas as code, which are evaluated on various optimization problems. singh2025mechanistic applies BO to enhance LLM’s internal behavior to boost the model’s zero-shot and few-shot performance. In contrast to these papers, we focus on the prompt optimization problem that arises when formulating LLM-based self-improving AI system design as a Bayesian Optimization problem. In line with this focus, schneider2025hyperbandbased and kong2025metapromptoptimizationllmbasedsequential formulate prompt optimization among the fixed set of candidate prompts as Bayesian optimization. While schneider2025hyperbandbased utilizes ideas from adversarial bandits, kong2025metapromptoptimizationllmbasedsequential feeds each prompt text into an embedding model, and the GP surrogate operates on those fixed embedding features to guide selection. However, unlike these papers, we actively generate new candidate prompts as the process proceeds, rather than pre-specifying a fixed set of candidate prompts.
@@
Next, our paper relates to the literature on efficient ad evaluation in digital marketing using bandit methods. Here, the goal is to iteratively improving the ads served given a fixed set of candidate ads. schwartz2017customer adopts Thompson Sampling with a hierarchical general linear model for a real display advertising campaign for a financial-services firm. geng2020online partitions JD.com’s platform users into disjoint sub-populations and applies a contextual Thompson sampling algorithm to maximize the expected advertiser payoff. aramayo2023multiarmed developed a contextual Thompson sampling algorithm to dynamically determine the list of house ads to display on the homepage of an electronic retailer to maximize the accumulated click-through rate of the ads. ba2022advertising used parametric Thompson Sampling, where a parametric structure links features (media, audience attributes) to conversion likelihood to optimally allocate ad-media and target-audience combinations in high-dimensional settings with low success rates. The main difference between these bandit-based papers and ours is the existence of a pre-defined set of candidate arms. Our paper focuses on creating new solutions by analyzing previous results. On the other hand, bandit methods aim to balance exploration of a pre-defined set of candidate ads optimally. The scope of our problem is also quite different. Our self-improving AI iterates over three steps–generation, evaluation, and analysis–whereas, bandits focus on efficiently executing the evaluation step. Indeed, one could employ a bandit method, such as best-arm-identification, in our evaluation step (instead of using A/B tests).
Finally, our paper relates to the literature on ad content. Researchers have studied many different features of ad creatives and their effect on consumer response, e.g., creativity (rosengren2020meta), discount information (biswas_2025), emotionality of content (macinnis_etal_2002), and specific text phrases (rutz_etal_2017). However, a limitation of this literature is that each study is designed to test one specific ad feature in one specific advertising format/context (e.g., TV ads, search ads), and there is not much consensus on how various features and estimated effects should be combined to optimize the ad creative and whether these effects can be ported to other formats and audiences. In this paper, we move away from manually manipulating individual ad features and instead treat the ad-creative optimization problem as a high-dimensional optimization problem in the prompt-space; and our self-improving AI framework can be used to automate and optimize ad-creative generation at scale for any given combination product, ad-format, and audience.
3 Problem Definition
@@
Let $\Phi$ denote a generative AI model (or more generally, an AI system) that, given a prompt $\pi\in\Pi$, produces a candidate solution $\Phi(\pi)$. [^2] LLM-based AI systems can also be modified by tuning the LLM’s internal parameters using LLM fine-tuning methods such as GRPO (liu2024deepseek). In this paper, we focus on prompt optimization, which fixes those internal parameters throughout. This candidate solution $\Phi(\pi)$, in conjunction with some environmental variables $x\in\mathcal{X}$ produces an output $y(\Phi(\pi),x)$, where $y$ is drawn from a distribution over an output space $\mathcal{Y}$. For example, in the ad creative optimization example from $\S$1, $\Phi$ denotes an image generation AI model (e.g., Gemini Flash), $\pi$ is an image generation prompt that reflects what the marketer wants the ad to depict, $\Phi(\pi)$ is the ad creative generated by prompt $\pi$, $x\in\mathcal{X}$ denotes a user on the ad platform and/orother contextual features that affect the ad’s effectiveness (e.g., seasonality or time of day), and a realization of $y(\Phi(\pi),x)$ represents that user/context $x$’s reaction (e.g., clicks, conversions, spend) to the ad $\Phi(\pi)$.
Throughout, we allow for the distribution over $\mathcal{X}$ to depend on the prompt $\pi$. This is important for practical applications where such dependencies are common; for example, in digital advertising platforms, recommendation systems often determine which users are shown the ad $\Phi(\pi)$ based on the ad content, which depends on $\pi$. To capture such dependencies, we denote the input distribution over $\mathcal{X}$ by $\mathcal{D}_{\mathcal{X}}(\pi)$. In a fully randomized test, the distribution over $\mathcal{X}$ is independent of $\pi$, and we can denote the distribution of $\mathcal{X}$ as $\mathcal{D}_{\mathcal{X}}$.
@@
To evaluate the output, we use a scoring rule: a function $r:\mathcal{Y}\to[0,1]$ that assigns a scalar effectiveness score to each $y\in\mathcal{Y}$. In the digital marketing example, $r$ is a marketer-defined function that maps observed user responses (clicks, conversions, spend) to a normalized scalar effectiveness score. Then, our optimization objective is to discover the prompt $\pi^{*}$ that maximizes the objective function:
@@
$J$ cannot be optimized by directly computing gradients, because the input distribution $\mathcal{D}_{\mathcal{X}}(\pi)$ can shift with $\pi$; even when $J(\pi)$ is differentiable with respect to $\pi$ and $y$’s randomness is not dependent on $\pi$, we have:
@@
where $d_{\pi}$ is the density of $\mathcal{D}_{\mathcal{X}}(\pi)$ (williams1992simple). Since $d_{\pi}$ and $\nabla_{\pi}\log d_{\pi}(x)$ are unobservable, we cannot compute gradients from observed rewards. This motivates the use of an external LLM that can analyze prior data and propose directions for improvement without explicitly computing gradients, leading to self-improving AI algorithms we describe below that iteratively optimize for JJ. [^3] In fully randomized experiments, the distribution term drops out and only the model term remains.
Self-improving AI algorithm:
Given a scoring rule $r$, a self-improving AI algorithm $\mathcal{A}$ optimizes $J(\pi):=\mathbb{E}[r(y(\Phi(\pi),x))]$ by iteratively analyzing the previous history, proposing a new prompt, and observing its noisy score. Formally, starting from the initial prompt $\pi_{0}$, a self-improving AI algorithm proceeds in discrete optimization iterations, indexed by $t=1,2,\ldots$; at the beginning of iteration $t$, the algorithm has access to the history of past prompts and their observed scores, then selects a new prompt $\pi_{t}$, deploys the corresponding candidate solution $\Phi(\pi_{t})$, and receives a score feedback on its performance. We define the score $s_{t}$ of the prompt $\pi_{t}$ as the empirical average reward at $t$, i.e.,
@@
where $L$ denotes the number of input samples $x_{t}^{(l)}$ we evaluate for each $t$. Note that we compare the performance of self-improving algorithms in terms of the progress of $s_{t}$ across iterations $t$.
@@
Fix a generative model $\Phi$ and a scoring rule $r$. For each iteration $t\geq 1$, let the history of past evaluations be $H_{t-1}\;:=\;\{(\pi_{\tau},\Phi(\pi_{\tau}),s_{\tau})\}_{\tau=0}^{t-1}$, where $s_{\tau}$ is the empirically evaluated score of $\pi_{\tau}$, defined as
@@
$\pi_{t}=\mathcal{A}_{t}(H_{t-1}).$
@@
    Analysis: Analyze the previous history $H_{t-1}$.
@@
    Generation: Propose the next prompts $\pi_{t}$ based on the analysis.